Parth (parthkohli):

Not as easy at it looks. u = sin(x) doesn't work.\[\int {\sin^2 x dx}\]

OpenStudy (lgbasallote):

use trig formulas \[\sin^2 x = \frac{1 - \cos (2x)}2\]

OpenStudy (lgbasallote):

if i recall right, that is the formula for sin^2 x

Parth (parthkohli):

\[\sin^2x = {1 - \cos ^2x}\]\[\implies{ {1 - {\cos(2x }) \over 2}}\]

Parth (parthkohli):

Yeah, right.

OpenStudy (lgbasallote):

so now integrate that

OpenStudy (anonymous):

integration by parts and the reduction formulae would work as well!

Parth (parthkohli):

@Omniscience can you explain?

OpenStudy (lgbasallote):

i doubt IBP would work

OpenStudy (anonymous):

It does!

OpenStudy (lgbasallote):

hmm yeah it does

Parth (parthkohli):

@lgbasallote I see what you did there, but let's see what Omni says :)

OpenStudy (anonymous):

for the reduction formulae google it; im too lazy to derive it right now :P

Parth (parthkohli):

Okay, well :)

OpenStudy (lgbasallote):

"for integration by parts you can do that on your own" ~Omni

OpenStudy (anonymous):

Ah yes, the famous "this part is left as an exercise for the reader." Also lazy, I just typed it into maxima and got \[\frac{x-\frac{sin(2x)}{2}}{2}\]

Parth (parthkohli):

I got my answer using lgb's method, but haha I couldn't get reduction formulae from Wikipedia!

OpenStudy (anonymous):

if i have time; i will derive it..if its not googeable

OpenStudy (anonymous):

wolfram alpha is the math wikipedia

OpenStudy (anonymous):

anyway; just use the trig method, easier

Parth (parthkohli):

Ho! Ho! Ho! I found an appropriate method somewhere ;) by the way, it's the same as LGB's.

OpenStudy (shubhamsrg):

you learnt complex nos ? you can use that too here..